# LAM: Large Avatar Model for One-shot Animatable Gaussian Head

### v1

@@ 3.3 New Expression Animation
$F(\vec{\beta},\vec{\theta},\vec{\phi})=\mathcal{S}(T_{P}(\vec{\beta},\vec{\theta},\vec{\phi}),\textbf{J}(\vec{\beta}),\vec{\theta},\mathcal{W}),$ (3)
@@
$T_{P}(\vec{\beta},\vec{\theta},\vec{\phi})=\bar{T}+B_{S}(\vec{\beta};\mathcal{S})+B_{P}(\vec{\theta};\mathcal{P})+B_{E}(\vec{\phi};\mathcal{E}),$ (4)
@@
$T_{G}(\vec{\theta},\vec{\phi})=\bar{G}+B_{P}(\vec{\theta};\mathcal{P})+B_{E}(\vec{\phi};\mathcal{E}),$ (5)
@@
$F_{G}(\vec{\theta},\vec{\phi})=\mathcal{S}(T_{G}(\vec{\theta},\vec{\phi}),\bar{J},\vec{\theta},\mathcal{W}).$ (7)
@@
Figure 2: Text to animatable Gaussian avatar generation and editing pipelines.
3.4 Text to Animatable Gaussian Avatar Generation and Editing.
Our approach takes full advantage of the shape information and animation attributes provided by the FLAME model and the robust capabilities of the DinoV2 image feature extractor, which is trained on extensive image datasets. Furthermore, after training on large-scale datasets, the scaling capability of the designed Transformer-based architecture allows our framework to adapt effectively to various input styles, including those derived from generated images. Building on this foundation, we have designed a streamlined pipeline for converting text prompts into animatable Gaussian avatars, along with an efficient style editing process, as illustrated in Fig. 2. To generate target portrait images based on textual prompts, we employ established text-to-image generation frameworks, such as Stable Diffusion [52]. The resulting images are then processed by our system to create animatable Gaussian avatars through a feed-forward mechanism. In addition, our framework is equipped to facilitate style editing; we can take an input image and utilize an image-to-image translation framework to modify the portrait’s style, such as changing the age or transforming it into a cartoon representation. This edited image is subsequently fed into our model to produce animatable Gaussian avatars that reflect the desired styles.
3.5 Deployment to Traditional Rendering Pipeline on Various Platforms.
@@ 3.6 Optimization and Regularization
$\mathcal{L}_{rgb}=\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{lpips}.$ (8)
@@
$\mathcal{L}_{o}=\mathcal{L}_{2}(O,\epsilon),$ (9)
@@
$\mathcal{L}=\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{lpips}+\lambda_{3}\mathcal{L}_{mask}+\lambda_{4}\mathcal{L}_{o}.$ (10)
Table 1: Quantitative results on the VFHQ dataset.
@@
StyleHeat [71] 19.95 0.726 0.211 0.537 0.199 0.385 7.659 0.407 0.279 0.551
ROME [33] 19.96 0.786 0.192 0.701 0.138 0.186 4.986 0.530 0.259 0.277
OTAvatar [43] 17.65 0.563 0.294 0.465 0.234 0.545 18.19 0.364 0.324 0.678
HideNeRF [38] 19.79 0.768 0.180 0.787 0.143 0.361 7.254 0.514 0.277 0.527
GOHA [39] 20.15 0.770 0.149 0.664 0.176 0.173 6.272 0.518 0.274 0.261
CVTHead [40] 18.43 0.706 0.317 0.504 0.186 0.224 5.678 0.374 0.261 0.311
GPAvatar [9] 21.04 0.807 0.150 0.772 0.132 0.189 4.226 0.564 0.255 0.328
Real3DPortrait [69] 20.88 0.780 0.154 0.801 0.150 0.268 5.971 0.663 0.296 0.411
Portrait4D [13] 20.35 0.741 0.191 0.765 0.144 0.205 4.854 0.596 0.286 0.258
Portrait4D-v2 [14] 21.34 0.791 0.144 0.803 0.117 0.187 3.749 0.656 0.268 0.273
GAGAvatar [7] 21.83 0.818 0.122 0.816 0.111 0.135 3.349 0.633 0.253 0.247
@@
Table 2: Quantitative results on the HDTF dataset.
@@
StyleHeat [71] 21.41 0.785 0.155 0.657 0.158 0.162 4.585 0.632 0.271 0.239
ROME [33] 20.51 0.803 0.145 0.738 0.133 0.123 4.763 0.726 0.268 0.191
OTAvatar [43] 20.52 0.696 0.166 0.662 0.180 0.170 8.295 0.643 0.292 0.222
HideNeRF [38] 21.08 0.811 0.117 0.858 0.120 0.247 5.837 0.843 0.276 0.288
GOHA [39] 21.31 0.807 0.113 0.725 0.162 0.117 6.332 0.735 0.277 0.136
CVTHead [40] 20.08 0.762 0.179 0.608 0.169 0.138 4.585 0.591 0.242 0.203
GPAvatar [9] 23.06 0.855 0.104 0.855 0.114 0.135 3.293 0.842 0.268 0.219
Real3DPortrait [69] 22.82 0.835 0.103 0.851 0.138 0.137 4.640 0.903 0.299 0.238
Portrait4D [13] 20.81 0.786 0.137 0.810 0.134 0.131 4.151 0.793 0.291 0.240
Portrait4D-v2 [14] 22.87 0.860 0.105 0.860 0.111 0.111 3.292 0.857 0.262 0.183
GAGAvatar [7] 23.13 0.863 0.103 0.862 0.110 0.111 2.985 0.851 0.231 0.181
@@
Table 3: Running time of reenactment and rendering measured in FPS. All results exclude the time for avatar reconstruction and driving parameters estimation that can be calculated in advance. The results are averaged over 100 frames.
@@
4 Experiments
4.1 Experiments Setting
@@
We utilize the VFHQ dataset [64] for training our model, which consists of video clips from a variety of interview scenarios. The dataset comprises 15,204 video clips with 3M frames. For each frame extracted from the videos, we detect the face region, enlarge the bounding box to crop out the region of interest, and resize the cropped images to 512×512 pixels for uniformity following GAGAvatar [7]. To enhance the dataset’s usability, we implement tracking for camera poses and FLAME parameters as in GaussianAvatar [50]. Following the methodology established by [8], we also conducted background removal to isolate the subjects in the frames. For evaluation purposes, we employ sampled frames from the original test split of the VFHQ dataset following previous works [8, 7]. In our setup, the first frame of each video is designated as the source image, while the subsequent frames act as driving and target images for the reenactment process. Additionally, we conduct evaluations using the HDTF dataset [77] adhering to the test split established by prior works [42]. This evaluation includes a selection of 19 video clips, complementing our analysis and contributing to a robust assessment of our model’s performance across different datasets.
@@
In our evaluation of single-image animatable Gaussian head avatar generation, we focus on self and cross-identity reenactment performance metrics. For self-reenactment, when ground truth images are available, we assess the quality of the synthesized images using three quantitative measures: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). These metrics allow us to compare the synthetic images against the ground truth effectively. To evaluate identity similarity (CSIM), we compute the cosine distance of face recognition features, building on the work of [11]. For assessing expression and pose fidelity, we employ the Average Expression Distance (AED) and Average Pose Distance (APD) as determined by a 3D Morphable Model (3DMM) estimator [12]. Additionally, we measure the Average Keypoint Distance (AKD) using a facial landmark detector as proposed by [5]. These metrics provide insights into the accuracy of the driving control in our animations. In cross-identity reenactment, where ground truth images are unavailable, we rely on CSIM, AED, and APD for evaluation. These metrics are in line with methodologies used in prior research [7], allowing for a coherent comparison across different approaches.
4.2 Main Results
@@
Table 4: Effect of different design choices on the VFHQ dataset.
@@ 4.3 More Applications
Fig. 4 shows the effective results of applying our framework for style editing of animatable Gassian avatars. Unlike previous 3D editing frameworks [24, 25, 62] that require iterative training on multi-view images for stylization, our framework can edit different styles of the 3D Gaussian avatar efficiently utilizing a 2D editing prior models to edit the avatar in the 2D image and then lift it to 3D Gaussian space, enabling efficient and user-friendly editing of 3D assets.
4.4 Ablation Studies
@@
We also compare widely used triplane-based query representation with our point-based query. Specifically, we replace the learnable query point features $F_{P}$ with flattened learnable triplane features for cross-attention with the image feature. Each sampled point then fetch its corresponding features from the triplane by projection for Gaussian attributes decoding. As shown in Table 4, our point-based framework (LAM-80K) gets better results compared with triplane-based (Tri.).
@@
We compare our canonical space generation with directly generating Gaussian avatar in the reference expression and pose (PE in Table 4) and then reenact it to the driven pose and expression. The canonical space simplifies the problem and gets better results.
@@
Compared with only applying cross-attention on the painted image feature as in [46] (denoted as Paint in Table 4), directly performing cross-attention on the image features gets quite better results.
5 Conclusions and Limitations
@@
Limitations. We utilize FLAME parameters to animate our reconstructed 3D Gaussian avatar. It fails to reproduce expressions that FLAME cannot model, i.e. sticking out the tongue, since FLAME does not model the tongue blendshapes. Also, since we utilize algorithms to estimate FLAME from video, the inaccurate estimated FLAME parameters will infect the results.
Figure 3: Visualization of cross-reenacted results on the generated image with text prompts.
Figure 4: Visualization of cross-reenacted results on the stylized images.
Figure 5: Visualization of cross-reenacted results on the VFHQ dataset.
@@ Appendix A More results
Figure 6: Visualization of the reenactment and rendering of our reconstructed Gaussian avatar on various computing platforms, including mobile phones.
Figure 7: Visualization of self-reenacted results on the VFHQ dataset.

### v2

@@ 3.3. New Expression Animation
(3) $F(\vec{\beta},\vec{\theta},\vec{\phi})=\mathcal{S}(T_{P}(\vec{\beta},\vec{\theta},\vec{\phi}),\textbf{J}(\vec{\beta}),\vec{\theta},\mathcal{W}),$
@@
(4) $T_{P}(\vec{\beta},\vec{\theta},\vec{\phi})=\bar{T}+B_{S}(\vec{\beta};\mathcal{S})+B_{P}(\vec{\theta};\mathcal{P})+B_{E}(\vec{\phi};\mathcal{E}),$
@@
(5) $T_{G}(\vec{\theta},\vec{\phi})=\bar{G}+B_{P}(\vec{\theta};\mathcal{P})+B_{E}(\vec{\phi};\mathcal{E}),$
@@
(7) $F_{G}(\vec{\theta},\vec{\phi})=\mathcal{S}(T_{G}(\vec{\theta},\vec{\phi}),\bar{J},\vec{\theta},\mathcal{W}).$
@@
Figure 3. Text to animatable Gaussian avatar generation and editing pipelines.
3.4. Text to Animatable Gaussian Avatar Generation and Editing.
Our approach takes full advantage of the shape information and animation attributes provided by the FLAME model and the robust capabilities of the DinoV2 image feature extractor, which is trained on extensive image datasets. Furthermore, after training on large-scale datasets, the scaling capability of the designed Transformer-based architecture allows our framework to adapt effectively to various input styles, including those derived from generated images. Building on this foundation, we have designed a streamlined pipeline for converting text prompts into animatable Gaussian avatars, along with an efficient style editing process, as illustrated in Fig. 3. To generate target portrait images based on textual prompts, we employ established text-to-image generation frameworks, such as Stable Diffusion (Rombach et al., 2022). The resulting images are then processed by our system to create animatable Gaussian avatars through a feed-forward mechanism. In addition, our framework is equipped to facilitate style editing; we can take an input image and utilize an image-to-image translation framework to modify the portrait’s style, such as changing the age or transforming it into a cartoon representation. This edited image is subsequently fed into our model to produce animatable Gaussian avatars that reflect the desired styles.
3.5. Deployment to Traditional Rendering Pipeline on Various Platforms.
@@ 3.6. Optimization and Regularization
(8) $\mathcal{L}_{rgb}=\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{lpips}.$
@@
(9) $\mathcal{L}_{o}=\mathcal{L}_{2}(O,\epsilon),$
@@
(10) $\mathcal{L}=\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{lpips}+\lambda_{3}\mathcal{L}_{mask}+\lambda_{4}\mathcal{L}_{o}.$
Table 1. Quantitative results on the VFHQ dataset.
@@
StyleHeat (Yin et al., 2022b) 19.95 0.726 0.211 0.537 0.199 0.385 7.659 0.407 0.279 0.551
ROME (Khakhulin et al., 2022b) 19.96 0.786 0.192 0.701 0.138 0.186 4.986 0.530 0.259 0.277
OTAvatar (Ma et al., 2023c) 17.65 0.563 0.294 0.465 0.234 0.545 18.19 0.364 0.324 0.678
HideNeRF (Li et al., 2023c) 19.79 0.768 0.180 0.787 0.143 0.361 7.254 0.514 0.277 0.527
GOHA (Li et al., 2023a) 20.15 0.770 0.149 0.664 0.176 0.173 6.272 0.518 0.274 0.261
CVTHead (Ma et al., 2024) 18.43 0.706 0.317 0.504 0.186 0.224 5.678 0.374 0.261 0.311
GPAvatar (Chu et al., 2024b) 21.04 0.807 0.150 0.772 0.132 0.189 4.226 0.564 0.255 0.328
Real3DPortrait (Ye et al., 2024) 20.88 0.780 0.154 0.801 0.150 0.268 5.971 0.663 0.296 0.411
Portrait4D (Deng et al., 2024b) 20.35 0.741 0.191 0.765 0.144 0.205 4.854 0.596 0.286 0.258
Portrait4D-v2 (Deng et al., 2024a) 21.34 0.791 0.144 0.803 0.117 0.187 3.749 0.656 0.268 0.273
GAGAvatar (Chu and Harada, 2024) 21.83 0.818 0.122 0.816 0.111 0.135 3.349 0.633 0.253 0.247
@@
Table 2. Quantitative results on the HDTF dataset.
@@
StyleHeat (Yin et al., 2022b) 21.41 0.785 0.155 0.657 0.158 0.162 4.585 0.632 0.271 0.239
ROME (Khakhulin et al., 2022b) 20.51 0.803 0.145 0.738 0.133 0.123 4.763 0.726 0.268 0.191
OTAvatar (Ma et al., 2023c) 20.52 0.696 0.166 0.662 0.180 0.170 8.295 0.643 0.292 0.222
HideNeRF (Li et al., 2023c) 21.08 0.811 0.117 0.858 0.120 0.247 5.837 0.843 0.276 0.288
GOHA (Li et al., 2023a) 21.31 0.807 0.113 0.725 0.162 0.117 6.332 0.735 0.277 0.136
CVTHead (Ma et al., 2024) 20.08 0.762 0.179 0.608 0.169 0.138 4.585 0.591 0.242 0.203
GPAvatar (Chu et al., 2024b) 23.06 0.855 0.104 0.855 0.114 0.135 3.293 0.842 0.268 0.219
Real3DPortrait (Ye et al., 2024) 22.82 0.835 0.103 0.851 0.138 0.137 4.640 0.903 0.299 0.238
Portrait4D (Deng et al., 2024b) 20.81 0.786 0.137 0.810 0.134 0.131 4.151 0.793 0.291 0.240
Portrait4D-v2 (Deng et al., 2024a) 22.87 0.860 0.105 0.860 0.111 0.111 3.292 0.857 0.262 0.183
GAGAvatar (Chu and Harada, 2024) 23.13 0.863 0.103 0.862 0.110 0.111 2.985 0.851 0.231 0.181
@@
Table 3. Running time of reenactment and rendering measured in FPS. All results exclude the time for avatar reconstruction and driving parameters estimation that can be calculated in advance. The results are averaged over 100 frames.
@@
4. Experiments
4.1. Experiments Setting
@@
We utilize the VFHQ dataset (Xie et al., 2022) for training our model, which consists of video clips from a variety of interview scenarios. The dataset comprises 15,204 video clips with 3M frames. For each frame extracted from the videos, we detect the face region, enlarge the bounding box to crop out the region of interest, and resize the cropped images to 512×512 pixels for uniformity following GAGAvatar (Chu and Harada, 2024). To enhance the dataset’s usability, we implement tracking for camera poses and FLAME parameters as in GaussianAvatar (Qian et al., 2024). Following the methodology established by (Chu et al., 2024a), we also conducted background removal to isolate the subjects in the frames. For evaluation purposes, we employ sampled frames from the original test split of the VFHQ dataset following previous works (Chu et al., 2024a; Chu and Harada, 2024). In our setup, the first frame of each video is designated as the source image, while the subsequent frames act as driving and target images for the reenactment process. Additionally, we conduct evaluations using the HDTF dataset (Zhang et al., 2021) adhering to the test split established by prior works (Ma et al., 2023b). This evaluation includes a selection of 19 video clips, complementing our analysis and contributing to a robust assessment of our model’s performance across different datasets.
@@
In our evaluation of single-image animatable Gaussian head avatar generation, we focus on self and cross-identity reenactment performance metrics. For self-reenactment, when ground truth images are available, we assess the quality of the synthesized images using three quantitative measures: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). These metrics allow us to compare the synthetic images against the ground truth effectively. To evaluate identity similarity (CSIM), we compute the cosine distance of face recognition features, building on the work of (Deng et al., 2022). For assessing expression and pose fidelity, we employ the Average Expression Distance (AED) and Average Pose Distance (APD) as determined by a 3D Morphable Model (3DMM) estimator (Deng et al., 2019). Additionally, we measure the Average Keypoint Distance (AKD) using a facial landmark detector as proposed by (Bulat and Tzimiropoulos, 2017). These metrics provide insights into the accuracy of the driving control in our animations. In cross-identity reenactment, where ground truth images are unavailable, we rely on CSIM, AED, and APD for evaluation. These metrics are in line with methodologies used in prior research (Chu and Harada, 2024), allowing for a coherent comparison across different approaches.
4.2. Main Results
@@
Table 4. Effect of different design choices on the VFHQ dataset.
@@ 4.3. More Applications
Fig. 5 shows the effective results of applying our framework for style editing of animatable Gassian avatars. Unlike previous 3D editing frameworks (Haque et al., 2023; He et al., 2024; Wang et al., 2024) that require iterative training on multi-view images for stylization, our framework can edit different styles of the 3D Gaussian avatar efficiently utilizing a 2D editing prior models to edit the avatar in the 2D image and then lift it to 3D Gaussian space, enabling efficient and user-friendly editing of 3D assets.
4.4. Ablation Studies
@@
We also compare widely used triplane-based query representation with our point-based query. Specifically, we replace the learnable query point features $F_{P}$ with flattened learnable triplane features for cross-attention with the image feature. Each sampled point then fetch its corresponding features from the triplane by projection for Gaussian attributes decoding. As shown in Table 4, with the same number of points, our point-based framework (LAM-80K) gets better results compared with triplane-based (Tri.).
@@
We compare our canonical space generation with directly generating Gaussian avatar in the reference expression and pose (PE in Table 4) and then reenact it to the driven pose and expression. The canonical space simplifies the problem and gets better results. We use the same number of points (80K) for fair comparisons.
@@
Compared with only applying cross-attention on the painted image feature as in (Pan et al., 2023) (denoted as Paint in Table 4), directly performing cross-attention on the image features gets quite better results. We use the same number of points (80K) for fair comparisons.
5. Conclusions and Limitations
@@
Limitations.
We utilize FLAME parameters to animate our reconstructed 3D Gaussian avatar. It fails to reproduce expressions that FLAME cannot model. For example, LAM is not able to model the tongue movement since FLAME does not model the tongue blend-shapes. Since we eliminate the 2D post-processing network for efficient animation and rendering, some expression-dependent details like dynamic wrinkles cannot be fully modeled. The limitation of FLAME expressiveness and the challenge of single input image also limits the expression neutralization capability. Also, since we utilize algorithms to estimate FLAME from video, the inaccurate estimated FLAME parameters will infect the results.
@@ Appendix A More results
Figure 7. Visualization of the reenactment and rendering of our reconstructed Gaussian avatar on various computing platforms, including mobile phones.
[…]
Figure 8. Visualization of self-reenacted results on the VFHQ dataset.
[…]